OpenStudy (anonymous):

On a trip from Williamsburg, Virginia to Orlando, Florida, a family wants to travel at least 250 miles in the first 5 hours of driving. What should their average speed be in order to meet their goal?

OpenStudy (anonymous):

V = d/ t = 250/5 = 50 mi/hr

OpenStudy (anonymous):

i dont understand what your trying to say?

OpenStudy (anonymous):

Do you know the formula to find velocity/ speed?

OpenStudy (anonymous):

no , but is the answer 50 ?

OpenStudy (anonymous):

Yes, the average speed is 50

OpenStudy (anonymous):

All you do is apply the formula!